### v1

# Risk-Aware Real-Time Task Allocation for

Abstract
This paper addresses the control synthesis of heterogeneous stochastic linear multi-agent systems with real-time allocation of signal temporal logic (STL) specifications. Based on previous work, we decompose specifications into sub-specifications on the individual agent level. To leverage the efficiency of task allocation, a heuristic filter evaluates potential task allocation based on STL robustness. Subsequently, an auctioning algorithm determines the definite allocation of specifications. Finally, a control strategy is synthesized for each agent-specification pair using tube-based Model Predictive Control (MPC), ensuring provable probabilistic satisfaction. We demonstrate the efficacy of the proposed methods using a multi-bus scenario that highlights a promising extension to autonomous driving applications like crossing an intersection.
I Introduction
Multi-agent systems are frequently employed in complex tasks that require close collaborations. Formal specifications, such as signal temporal logic (STL), are increasingly utilized to define complex collaborative task objectives [1]. The collaborative control of multi-agent systems satisfying STL specifications has been studied using mixed-integer programming [2], funnel functions [3], and control barrier functions [4]. Nevertheless, a critical challenge for the practical implementation of these methods is their reliability in uncertain and dynamic environments. For systems subject to stochastic disturbances, it is important to evaluate the impact of these disturbances on the task accomplishment of individual agents, so that the associated risks are restricted [5]. When tasks are assigned at runtime [6], task allocation and control become challenging in the multi-agent framework.
Despite recent efforts, there exists a notable gap in the literature regarding the control of stochastic multi-agent systems under real-time allocated STL specifications with probabilistic guarantees. Such systems have significant practical value, exemplified by scenarios such as the two-vehicle intersection case in Figure 1. In such a scenario, each vehicle must navigate while considering its stochastic disturbances and avoiding collisions with its opponent. The complexity of the problem necessitates decomposing the overarching driving specification into two sub-specifications tailored to each vehicle. These sub-specifications, characterized by different time intervals, may require real-time assignment. Hence, this paper addresses the control synthesis of stochastic multi-agent systems under collaborative STL specifications assigned at runtime with probabilistic guarantees.
Figure 1: Illustration of decomposing tasks for autonomous driving.
A crucial aspect of addressing this challenging problem involves task assignment or allocation (TA), which aims to determine optimal agent-task pairs in a multi-agent system [7]. This process yields a linear programming problem  [8], efficiently solvable using auction algorithms [9]. Dynamically assigned tasks have been described using real-time STL specifications [10, 11]. TA necessitates decomposing specifications assigned to multiple collaborative agents into individual agent specifications, requiring the decomposition of STL specifications both among agents [12, 13] and across time [14]. With decomposed real-time specifications, control synthesis for stochastic multi-agent systems holds promise with the utilization of shrinking-horizon model predictive control (MPC) [15] to ensure bounded risks [16]. In previous work, probabilistic reachable tubes have been utilized to optimize risk bounds for both predefined specifications [17] and real-time specifications [6]. These methods underpin the risk-aware control synthesis of stochastic multi-agent systems with decomposed real-time specifications.
This paper proposes a novel approach to allocating STL specifications, involving subgroups of agents, and synthesizing control strategies for individual agents in real-time. By decomposing newly assigned specifications into agent-level tasks, we employ a heuristic approach based on STL robustness [18] to assess agent-specification pairs and attack task assignment via an auctioning scheme formulated as a mixed integer linear program. Subsequently, a tube-based model predictive control strategy is employed where an optimization is solved at every time step at agent level. The solution of this optimization ensures the satisfaction of running and newly allocated tasks with a desired risk level, while discarding tasks not adhering to the risk specification. We prove recursive feasibility of this MPC scheme, leading to an open-loop system with provable probabilistic guarantees and a closed-loop system for which guarantees are suggested.
II Preliminaries and Problem Statement
This section introduces the preliminaries, class of systems, temporal logics, and basic assumptions, culminating in a problem statement and an approach.
II-A Preliminaries
For a given probability measure $\mathbb{P}$ defined over Borel measurable space $(\mathbb{X},\mathcal{B}(\mathbb{X}))$, we denote the probability of an event $\mathcal{A}\in\mathcal{B}(\mathbb{X})$ as $\mathbb{P}(\mathcal{A})$. In this paper, we will work with Euclidean spaces and Borel measurability. Further details on measurability are omitted, and we refer the interested reader to [19]. Additionally, any half-space $H\subset\mathbb{R}^{n}$ is defined as $H:=\{x\in\mathbb{R}^{n}\mid g^{T}x\geq b\}$ with $g\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$. A polyhedron is the intersection of finitely many half-spaces, also denoted as $H:=\{x\in\mathbb{R}^{n}\mid Gx\geq b\}$ with $G\in\mathbb{R}^{q\times n}$ and $b\in\mathbb{R}^{q}$. A polytope is a bounded polyhedron. The 2-norm is defined by $||x||=\sqrt{x^{T}x}$ with $x\in\mathbb{R}^{n}$. The identity matrix is denoted by $I_{n}\in\mathbb{R}^{n\times n}$. The vector of all elements one is denoted by $\bm{1}_{n}\in\mathbb{R}^{n}$, but if the context is clear, it will instead be denoted by $\bm{1}$. The zero vector is denoted by $\bm{0}_{n}\in\mathbb{R}^{n}$, but if the context is clear, it will instead be denoted by $\bm{0}$. For any finite set $I$, the cardinality is denoted by $|I|$. $\mathrm{Proj}_{X}(Y)$ denotes the projection of elements in $Y$ onto $X$.
II-B Multi-agent Systems
In this paper, we consider a finite ordered set $\mathcal{I}$ of heterogeneous agents where each agent $i\in\mathcal{I}$ has linear time-invariant dynamics with additive noise, given by
$x^{i}(k+1)=A^{i}x^{i}(k)+B^{i}u^{i}(k)+w^{i}(k),$ (1)
[…]
For each agent $i\in\mathcal{I}$, we assume there is a local controller. We define the local controllers as a sequence of policies
$\bm{f}^{i}:=\{f^{i}_{0},f^{i}_{1},\dots\},$
such that $f^{i}_{k}\!\!:\!\!\mathbb{H}^{i}_{k}\!\!\to\!\!\mathbb{U}^{i}$ maps the history of states and inputs to inputs. Here $\mathbb{H}^{i}_{k}\!:=\!(\mathbb{X}^{i}\times\mathbb{U}^{i})^{k}\times\mathbb{X}^{i}$ with elements $\eta^{i}(k)\!:=\!(x^{i}(0),u^{i}(0),\cdots,u^{i}(k-1),x^{i}(k))$. By implement-ing controller $\bm{f}^{i}$ upon the corresponding agent, we obtain the controlled form of (1) for which the control input satisfies the feedback law $u^{i}(k)=f^{i}_{k}(\eta^{i}(k))$ with $\eta^{i}(k)\in\mathbb{H}^{i}_{k}$. We indicate the input sequence of agent $i$ by $\bm{u}^{i}:=\{u^{i}(0),u^{i}(1),\dots\}$ and define its executions as sequences of states $\bm{x}^{i}:=\{x^{i}(0),x^{i}(1),\dots\}$ referred to as signals.
We define the suffix, segment and signal element of any signal $\bm{x}^{i}:=\{x^{i}(0),x^{i}(1),\ldots\}$, respectively, by $\bm{x}^{i}_{k}=\{x^{i}(k),x^{i}(k+1),\dots\}$, $\bm{x}^{i}_{[a,b]}:=\{x^{i}(a),\ldots,x^{i}(b)\}$ with $a\leq b$, and $\bm{x}^{i}(k):=\bm{x}^{i}_{[k,k]}$. We define the concatenation of $\bm{x}^{i}$, for all agents $i\in I\subseteq\mathcal{I}$, by a higher dimensional signal $\bm{x}^{I}$, where $\bm{x}^{I}(k):=[\bm{x}^{i_{1}}(k)^{T},\cdots,\bm{x}^{i_{l}}(k)^{T}]^{T}$, $i_{j}\in I$, and $l=|I|$ is the number of elements in $I$. Any signal $\bm{x}^{i}$ can be interpreted as a realization of the probability distribution induced by implementing controller $\bm{f}^{i}$, denoted by $\bm{x}^{i}\sim\mathbb{P}_{\bm{f}^{i}}$. Similar realizations exist for segments, i.e., $\bm{x}^{i}_{k}\sim\mathbb{P}_{\bm{f}^{i},k}$ and concatenations, i.e., $\bm{x}^{I}\sim\mathbb{P}_{\bm{f}^{I}}$, with $\bm{f}^{I}:=\{\bm{f}^{i}\mid i\in I\}$.
II-C Signal Temporal Logic & Probabilistic Satisfaction
To mathematically describe tasks, e.g., reaching a target within a given time frame, we consider specifications given by signal temporal logic. Here, we assume that all STL specifications adhere to the negation normal form (NNF), see [1]. This assumption does not restrict the overall framework since [21] proves that every STL specification can be rewritten into the negation normal form.
STL consists of predicates $\mu^{l}$ for subgroups of $l$ agents that are either true ($\top$) or false ($\bot$). Each predicate $\mu^{l}$ is described by a function $h:\mathbb{R}^{nl}\to\mathbb{R}^{q}$, $l\leq|\mathcal{I}|$, as follows
$\mu^{l}:=\begin{cases}\top\text{ if }\forall i\in\{1,\cdots,q\},\ h_{i}(x)\geq 0\\ \bot\text{ if }\exists i\in\{1,\cdots,q\},\ h_{i}(x)<0.\end{cases}$ (2)
Here $h_{i}:\mathbb{R}^{nl}\to\mathbb{R}$ indicates the $i$-th element of function $h$. We will assume that all predicate functions are affine functions of the form $h(x)=Gx-b$ with $G\in\mathbb{R}^{q\times nl}$, $b\in\mathbb{R}^{q}$ and $||g_{j}||=1$ for all $j\in\{1,\cdots,q\}$, where $g_{j}$ is the $j$-th row of $G$. Furthermore, we assume that $H_{\mu^{l}}:=\{x\in\mathbb{R}^{nl}\mid Gx\geq b\}$ describes a polytope. Let the set of all predicates be given by $\mathcal{P}$. The STL syntax will be given by
$\phi::=\top\mid\mu^{l}\mid\lnot\mu^{l}\mid\phi_{1}\wedge\phi_{2}\mid\phi_{1}\vee\phi_{2}\mid\square_{[a,b]}\phi\mid\phi_{1}\ U_{[a,b]}\ \phi_{2},$
where $\mu^{l}\in\mathcal{P}$, $\phi$, $\phi_{1}$ and $\phi_{2}$ are STL formula, $a,b\in\mathbb{N}$, and $a\leq b$. The semantics are given next, where $\bm{x}^{\mathcal{I}}_{k}\vDash\phi$ denotes the satisfaction of $\phi$ verified over the suffix of signal $\bm{x}^{\mathcal{I}}$.
Definition 1
The STL semantics, for a specification jointly assigned to a set of agents, are recursively given by:
$\bm{x}^{\mathcal{I}}_{k}\vDash\mu^{l}$ $\iff$ $\exists I\subseteq\mathcal{I}:\ |I|=l,\ h({\bm{x}^{I}(k)})\geq\bm{0}$
$\bm{x}^{\mathcal{I}}_{k}\vDash\lnot\mu^{l}$ $\iff$ $\nexists I\subseteq\mathcal{I}:\ |I|=l,\ h({\bm{x}^{I}(k)})\geq\bm{0}$
$\bm{x}^{\mathcal{I}}_{k}\vDash\phi_{1}\wedge\phi_{2}$ $\iff$ $\bm{x}^{\mathcal{I}}_{k}\vDash\phi_{1}\text{ and }\bm{x}^{\mathcal{I}}_{k}\vDash\phi_{2}$
$\bm{x}^{\mathcal{I}}_{k}\vDash\phi_{1}\vee\phi_{2}$ $\iff$ $\bm{x}^{\mathcal{I}}_{k}\vDash\phi_{1}\text{ or }\bm{x}^{\mathcal{I}}_{k}\vDash\phi_{2}$
$\bm{x}^{\mathcal{I}}_{k}\vDash\square_{[a,b]}\ \phi$ $\iff$ $\forall k^{\prime}\in[k+a,k+b]:\bm{x}^{\mathcal{I}}_{k^{\prime}}\vDash\phi$
$\bm{x}^{\mathcal{I}}_{k}\vDash\phi_{1}\ U_{[a,b]}\ \phi_{2}$ $\iff$ $\exists k_{1}\in[k+a,k+b]:\bm{x}^{\mathcal{I}}_{k_{1}}\vDash\phi_{2}$
[…]
Here, we introduced a slight abuse of the notation as $[a,b]$ is used to describe the set of all natural numbers within the real interval $[a,b]$. Additional operators can be derived such as the eventually-operator $\lozenge_{[a,b]}\phi:=\top U_{[a,b]}\phi$.
Since all agents behave stochastically, each STL specification can only be satisfied probabilistically. Should a specification be assigned at time $k$, the probability can be determined based on the state measurement $x^{\mathcal{I}}(k)$, the controller $\bm{f}^{\mathcal{I}}$, and the system dynamics (1). Accordingly, we consider the probability that suffix fragment $\bm{x}^{\mathcal{I}}_{k}$ satisfies specification $\phi$, given state $x^{\mathcal{I}}(k)$. We denote this by
$\mathbb{P}_{\bm{f}}(\phi,k):=\mathbb{P}_{\bm{f}^{\mathcal{I}},k}(\bm{x}^{\mathcal{I}}_{k}\vDash\phi\mid x^{\mathcal{I}}(k)).$ (3)
For each specification $\phi$, we assume that given is the maximal allowable risk $r_{\phi,\max}$ of not satisfying specification $\phi$. In the remainder, we will refer to this as the maximal risk. Note that for $|\mathcal{I}|=1$, the above definitions of STL and probability will become the standard ones as seen in [6].
II-D Problem Statement & Approach
We first consider an example (see Figure 2) to illustrate the problem and approach covered within this paper. Next, we give a problem description and a problem statement. Lastly, we elaborate upon the approach as seen in Figure 3.
Example 1. We examine a multi-robotic scenario in which robots are assigned joint tasks. Upon receiving a new task, only its description and the required number of robots are known. A procedure decides which group of robots, if any, will perform the task. Given the assumption that all robots’ behaviour is independent, new tasks are initially decomposed into sub-tasks at the individual level; that is, each robot performs part of the task without requiring information from other robots. Subsequently, the procedure prompts potential robot sub-task combinations to investigate an update to the individual control strategy without disrupting ongoing sub-tasks. Among all combinations, the optimal assignment is selected and all control strategies are updated accordingly.
Figure 2: Illustration of decomposing tasks to the individual level and combining tasks with robots. Blue are robots and red are tasks.
Problem Description. Assume each agent $i\in\mathcal{I}$ has dynamics (1) with noise distribution $\mathcal{Q}^{i}_{w}$ and state update $x^{i}(k)$, at each time step $k\in\mathbb{N}$. The objective is to update the suffix controller $\bm{f}^{\mathcal{I}}_{k}$ at each time step $k\in\mathbb{N}$. Here, any newly provided specification is either accepted or rejected based on the maximal risks of the newly provided specification and previously accepted specifications.
Problem Statement. Given a sequence of STL specifications $\bm{\phi}:=\{\phi_{1},\cdots,\phi_{t}\}$ assigned, respectively, at time instances $\{k_{1},\cdots,k_{t}\}$ with $k_{i}<k_{j}$ for $i<j$, with maximal risk $r_{\phi_{i},\max}$, develop a method for updating, at each time step $k\in\mathbb{N}$, suffix $\bm{f}^{\mathcal{I}}_{k}$ such that $\mathbb{P}_{\bm{f}}(\phi_{i},k_{i})\geq 1-r_{\phi_{i},\max}$ if $k_{i}\leq k$, where $\phi_{i}$ with $k_{i}=k$ is allowed to be rejected.
Approach. First, a newly provided STL specification and maximal risk are decomposed onto the single-agent level (Section IV-A). Subsequently, a filter determines which agent-specification pairs are favorable based on STL robustness and heuristics (Section IV-B). Afterwards, for each remaining agent-specification pair, a potential control strategy is determined together with the local risk value (Section III). Based on the local risk value for each feasible agent-specification pair, auctioning determines the definitive assignment (Section IV-C). Should no valid assignment exist, the specification will be rejected. Finally, the control strategy for each agent is updated.
Figure 3: Illustration of the developed approach at time $k$. (0) Determine for each agent-specification pair a control strategy and the local risk value. (1) Auctioning offers from agent-specification pairs. (2) Decision after auctioning. Red is covered in Section III and orange in Section IV.
III Risk-Aware Single-Agent Control Synthesis
Before the multi-agent case, this section considers the single-agent case. Consequently, every specification is on the single-agent level, and only a single agent-specification pair exists. Thus, decomposition of the specification, filtering of the pairs, and auctioning are omitted. Instead, this section will cover control synthesis and local risk value computation for any individual agent-specification pair.
The section will be structured as follows. First, we will consider decomposing the single-agent dynamics, define probabilistic reachable tubes and relate each specification to its maximal risk via (mostly) linear constraints. Second, we utilize the previously mentioned to formulate a tube-based MPC problem to synthesize a control strategy and determine the local risk value for any given agent-specification pair. This section may be regarded as a self-contained summary of our previous work on risk-aware MPC for single-agent systems [6]. We refer the reader to [6] for more details.
III-A Decomposing the Single-Agent Dynamics
In this paper, we decompose dynamics (1), given by
$x^{i}(k+1)=A^{i}x^{i}(k)+B^{i}u^{i}(k)+w^{i}(k),$
into a nominal and an error part [22]. The nominal dynamics, denoted as $z^{i}$, contain no stochasticity, and the (stochastic) error dynamics, denoted as $e^{i}$, are autonomous. This yields
$z^{i}(k+1)$ $=A^{i}z^{i}(k)+B^{i}v^{i}(k),$ (4a)
$e^{i}(k+1)$ $=A^{i}_{K}e^{i}(k)+w^{i}(k),$ (4b)
with $\mbox{with }x^{i}(k)$ $=z^{i}(k)+e^{i}(k),$ (4c)
$u^{i}(k)$ $=v^{i}(k)+K^{i}e^{i}(k).$ (4d)
Here $A^{i}_{K}=A^{i}+B^{i}K^{i}$ and $K^{i}$ is a stabilizing feedback gain meant to keep the error $e^{i}$ small. Throughout this section, we omit superscripts indicating agent indices from our notation.
III-B Probabilistic Reachable Tubes
We define probabilistic reachable sets similar to [22].
Definition 2 (Probabilistic Reachable Sets)
A probabili-stic reachable set (PRS) of dynamic probability level $p$ for error dynamics (4b), denoted by $\mathcal{R}^{p}$, is a set that satisfies
$e(0)=\bm{0}\Rightarrow\mathbb{P}(e(k)\in\mathcal{R}^{p})\geq p,\ \forall k\geq 0.$ (5)
Definition 2 states that a PRS of probability level $p$ will contain at any time instance $k$ the accumulating error $e(k)$ with at least probability $p$, if the initial error satisfies $e(0)=\bm{0}$. Multiple representations for probabilistic reachable sets exist, including the ellipsoidal representation. The ellipsoidal representation is obtained from the multivariable Chebyshev inequality; details are given in [22, 23]. For simplicity, we will make the following assumption.
Assumption 1
Distribution $\mathcal{Q}_{w}$ has zero mean.
The above assumption is not necessary to obtain an ellipsoidal representation but will simplify most of the computations. Under Assumption 1, the ellipsoidal representation is given by
$\mathcal{R}^{p}=\{x\in\mathbb{R}^{n}\mid x^{T}\Sigma_{\infty}^{-1}x\leq\tilde{p}\},$ (6)
where $\Sigma_{\infty}$ solves the Lyapunov equation $A_{K}\Sigma_{\infty}A_{K}^{T}+\text{var}(\mathcal{Q}_{w})=\Sigma_{\infty}$, $\tilde{p}=\frac{n}{1-p}$, and $p$ is the probability level. Since the variance of $\mathcal{Q}_{w}$ is assumed to be strictly positive definite and $A_{K}$ is stable, Lyapunov theory states that $\Sigma_{\infty}$ is strictly positive definite. In the sequel, we consider the transformed state and constraints for which $\Sigma_{\infty}=I_{n}$. These can be obtained by a simple coordinate transformation $y=\Sigma_{\infty}^{-\frac{1}{2}}x$. We call the transformed dynamics $y$ the normalized dynamics, and in the remainder, we maintain our notations as if the dynamics are normalized.
By defining a PRS $\mathcal{R}^{p_{k}}$ at each time instance $k\in\{0,\cdots,N\}$, we obtain a finite sequence of PRS called a probabilistic reachable tube (PRT) :=$\mathcal{R}:=\{\mathcal{R}^{p_{0}},\cdots,\mathcal{R}^{p_{N}}\}$. We define the risk level $r_{k}=1-p_{k}$ as an upper bound on the probability that $\mathcal{R}^{p_{k}}$ does not contain the error $e(k)$.
III-C Relating Specification and Maximal Risk
In this subsection, we aim to relate each STL specification $\phi$ to its maximal risk $r_{\phi,\max}$ via (mostly) linear constraints. To achieve this, we relate each STL specification $\phi$ assigned at time $k$, by means of a PRT, to the risk $r_{\phi}^{k}$ of failing specification $\phi$ assigned at time $k$. Subsequently, the desired relation can be derived by imposing that $r_{\phi}^{k}\leq r_{\phi,\max}$.
We first define the active horizon $\mathcal{H}_{\phi}^{k}$ of an STL formula $\phi$ as the set of all time instances needed to evaluate $\bm{x}_{k}\vDash\phi$ (see [6, Definition 4]). Here, it should be noted that any active horizon consists of only finite elements. Notice that the suffix $\bm{x}_{N+1}$ with $N=\max{\mathcal{H}_{\phi}^{k}}$, does not influence the satisfaction of $\bm{x}_{k}\vDash\phi$. Hence, in the remainder of this section, we consider only the prefix $\bm{x}_{[k,N]}:=\{x(k),\cdots,x(N)\}$ and note that $\mathbb{P}_{\bm{f},k}({\bm{x}}_{k}\vDash\phi\mid x(k))=\mathbb{P}_{\bm{f},k}({\bm{x}}_{[k,N]}\vDash\phi\mid x(k))$.
Given an STL specification $\phi$ assigned at time $k$, [1] explains that mixed integer linear constraints can represent any STL specification in negation normal form. As can be observed in the following equations, for each time instance $\tau\in\mathcal{H}_{\phi}^{k}$, these STL specifications define a linear constraint that is parametrized with binary variable $s$, as follows
$G(\tau)x(\tau)+MF(\tau)s$ $\geq b(M,\tau)+\rho(\tau)\bm{1},$ (7a)
$Cs\geq d,\ \epsilon\leq\rho(\tau)$ $\leq\frac{M}{2},\ \forall\tau\in\mathcal{H}_{\phi}^{k},$ (7b)
where $G(\tau)$, $E(\tau)$, $C$ and $b(M,\tau)$, $d$ are, respectively, matrices and vectors of appropriate sizes; $M$ is a large positive number; $\epsilon$ is a small positive number; and $\rho(\tau)>0\in\mathbb{R}$ is a positive real number.
We consider the following proposition to quantify risk $r_{\phi}^{k}$. The proof follows from [6, Theorem 5], which utilizes a PRT with risk levels $r(\tau)=\frac{n}{\rho(\tau)^{2}}$.
Proposition 3
Let specification $\phi$ be assigned at time $k$ to the system (4) with normalized dynamics. If the nominal trajectory $\{z(k),\cdots,z(N)\}$ with $z(k)=x(k)$ and nominal controls $\{v(k),\cdots,v(N-1)\}$ satisfies (7) with $\{\rho(k),\cdots,\rho(N)\}$ strictly positive, then
$\mathbb{P}_{\bm{f},k}({\bm{x}}_{k}\vDash\phi\mid x(k))\geq 1-r_{\phi}^{k},$
where risk $r_{\phi}^{k}$ is at most
$r_{\phi}^{k}=\sum_{\tau\in\mathcal{H}_{\phi}^{k}\backslash\{k\}}\frac{n}{\rho(\tau)^{2}}.$ (8)
A relation between specification $\phi$ assigned at time $k$ and risk $r_{\phi}^{k}$ follows from constraints (7) and (8).
III-D Control Synthesis and Local Risk Value
In this subsection, we compute the local risk value and synthesize a controller for a given agent-specification pair. We assume a new specification $\phi$ is provided at time $k$ together with maximal risk $r_{\phi,\max}$. For any specifications $\psi$, we denote with $k_{\psi}$ the time at which the specification was assigned. We denote with $\mathcal{P}^{k}$ the set of all previously accepted specifications $\psi$ with $k_{\psi}<k$, together with the newly assigned specification $\phi$, i.e., $\phi\in\mathcal{P}^{k}$. The objective is to synthesize a controller $\bm{f}$ such that $\varphi=\wedge_{\psi\in\mathcal{P}^{k}}\lozenge_{[k_{\psi},k_{\psi}]}\psi$ is satisfied in accordance to the maximal risk of each conjunction element, i.e., $\mathbb{P}_{\bm{f}}(\psi,k_{\psi})\geq 1-r_{\psi,\max},\ \forall\psi\in\mathcal{P}^{k}$.
We are interested in implementing a tube-based MPC algorithm that at each time instance $k\in[0,N]$ recomputes the optimal nominal trajectory $\bm{z}_{[k,N]}$, the optimal nominal input $\bm{v}_{[k,N-1]}$, and the optimal risk levels $\bm{r}_{[k,N]}$. For this, we assume given $x(k)$ together with $\bm{z}_{[0,k]}$, $\bm{v}_{[0,k-1]}$, $\bm{\rho}_{[0,k]}$ and $\bm{r}_{[0,k]}$ computed at time $k-1$. Furthermore, for all $\psi\in\mathcal{P}^{k}$, we assume $\mathcal{H}_{\psi}^{k_{\psi}}\subseteq[0,N]$ without lose of generality.
Let us consider the following tube-based MPC problem
$\min_{\Xi}\ J($ $\bm{z}_{[0,N]},\bm{v}_{[0,N-1]},\bm{r}_{[0,N]}),$ (9a)
[…]
$z(\tau+1)$ $=Az(\tau)+Bv(\tau),\forall\tau\in\{k,\cdots,N-1\},\!\!$ (9c)
$G(j)z(j)$ $+MF(j)s(k)\geq b(M,j)+\rho(j)\bm{1},$ (9d)
$Cs(k)$ $\geq d,\ \epsilon\leq\rho(j)\leq\textstyle\frac{M}{2},\ \forall j\in\{0,\cdots,N\},$ (9e)
$n$ $=\rho(j)^{2}r(j),\hskip 36.98866pt\forall j\in\{0,\cdots,N\},$ (9f)
$r_{\psi,\max}$ $\geq\textstyle\sum_{j\in\mathcal{H}_{\psi}^{k_{\psi}}\setminus\{k_{\psi}\}}r(j),\forall\psi\in\mathcal{P}^{k},$ (9g)
where $\Xi:=\{\bm{z}_{[k,N]},\bm{v}_{[k,N-1]},\bm{r}_{[k+1,N]},\bm{\rho}_{[k+1,N]},s(k)\}$, $J$ is either a linear or quadratic cost function, the constraints (9d)-(9e) are obtained from $\varphi=\wedge_{\psi\in\mathcal{P}^{k}}\lozenge_{[k_{\psi},k_{\psi}]}\psi$ and the constraints (9f)-(9g) are obtained from Prop. 3. From the above tube-based MPC, an (updated) control strategy is obtained from $\bm{v}_{[k,N-1]}$, and a local risk value $r_{\phi,mpc}$ is obtained from the right-hand side of (9g), i.e.,
$r_{\phi,mpc}=\sum_{j\in\mathcal{H}_{\phi}^{k}\setminus\{k\}}r(j).$ (10)
A solution to (9), by Prop. 3, ensures that for all $\psi\in\mathcal{P}^{k}$, $\mathbb{P}_{\bm{f}}(\psi,k_{\psi})\geq 1-r_{\psi,\max}$. Should there exist no solution to problem (9), the corresponding agent-specification pair is deemed infeasible. Consequently, in the single-agent case, we will then update the control strategy based only on measurement $x(k)$, as per constraint (9b). This can be achieved by considering the MPC problem (9) with $\mathcal{P}^{k}=\mathcal{P}^{k}\setminus\{\phi\}$. However, should also the measurement fail to produce a controller, the previously computed control strategy will be implemented instead.
IV Risk-Aware Multi-Agent Task Allocation
In this section, we consider the multi-agent case. More specifically, this section will primarily cover real-time task allocation for multi-agent systems. We refer to section III regarding control synthesis and local risk value computation for individual agent-specification pairs.
The section will be structured as follows. First, we will consider specification decomposition and heuristic agent-specification filtering. Second, we utilize auctioning, based on local risk values, to obtain a definitive agent-specification assignment. The section concludes with an implementation algorithm and theoretical analysis of the overall method.
IV-A Specification Decomposition
Let $\phi$ be a newly assigned STL specification with maximal risk $r_{\phi,\max}$, involving a distinct number of agents, $\nu\leq|\mathcal{I}|$. The objective is to decompose the specification $\phi$ into agent-level sub-specifications, $\phi^{j}$, $j\in\mathcal{I}$, and allocate individual maximal risks, $r_{\phi^{j},\max}$, in accordance to $r_{\phi,\max}$. First, we define $\mathrm{dec}(\phi,\nu)=\wedge_{j=1}^{\nu}\phi^{j}$ as the decomposition of $\phi$ into a conjunctive form, with each conjunct representing a sub-specification, $\phi^{j}$, the satisfaction of which depends on a single agent $j$, and $\nu$ being the number of agents needed to satisfy $\phi$. Second, we assign a maximal risk to $\phi^{j}$, following a union-bound argument, which requires $\sum_{j=1}^{\nu}r_{\phi^{j},\max}\geq r_{\phi,\max}$. In this paper, we select a uniform allocation of risks across sub-specifications, obtaining
$r_{\phi^{j},\max}=\frac{r_{\phi,\max}}{\nu},$ (11)
for all agents $j\in\{1,\ldots,\nu\}$.
The decomposition $\mathrm{dec}(\phi,\nu)$ acts as a transformation on $\phi$ returning a conjunctive formula, the satisfaction of which implies satisfaction of $\phi$. However, the converse is not necessarily true. There are a few methods [12, 13, 24] for splitting STL specifications into sub-specifications. In this paper, given a specification $\phi$, we convert Boolean and temporal operators to conjunctive forms based on the heuristics in [24] and decompose predicates following a similar approach to [12]. Therein, predicates are defined on general convex sets, whereas negated forms are not studied. Here, we focus on polytopic settings and allow negations. For our setting, we assume the following.
Assumption 2
Let $\phi$ be an STL specification assigned at time $k$. There exists decomposition $\mathrm{dec}(\phi,\nu)=\wedge_{j=1}^{\nu}\phi^{j}$, where $\phi^{j}$ indicates a task involving a single agent.
In view of existing decomposition approaches, where sub-specifications are assigned to subteams of agents, Assumption 2 is restrictive. However, in the setting of polytopic predicates, we can produce agent-level sub-specifications by projecting the associated predicates onto the agents’ state space. Before showing this, we define the following.
Definition 4
Let $H:=\{x\mid h(x)\geq\bm{0}\}$ be a polytope in $\mathbb{R}^{n}$. $B_{\mathrm{in}}(H):=\{x\mid\underline{b}_{j}^{\mathrm{in}}\leq x_{j}\leq\overline{b}_{j}^{\mathrm{in}},\;j=1,\ldots,n\}$ is an orthotope inscribed in $H$, i.e., $B_{\mathrm{in}}(H)\subseteq H$, and $B_{\mathrm{out}}(H):=\{x\mid\underline{b}_{j}^{\mathrm{out}}\leq x_{j}\leq\overline{b}_{j}^{\mathrm{out}},\;j=1,\ldots,n\}$ is an orthotope circumscribing $H$, i.e., $B_{\mathrm{out}}(H)\supseteq H$.
Lemma 5
Let $\mu:=(h(x)\geq\bm{0})$ be a predicate, $H:=\{x\mid h(x)\geq\bm{0}\}$ be a polytope, with $x=[x_{1}^{\intercal}\;\ldots\;x_{\nu}^{\intercal}]^{\intercal}$, $x_{i}\in\mathbb{R}^{n}$, $i\in\{1,\ldots,\nu\}$, and $B_{\mathrm{in}}(H)\subseteq H$ and $B_{\mathrm{out}}(H)\supseteq H$ be the associated orthotopes inscribed in and circumscribing $H$, respectively. Then, it holds that (i) $\mathrm{dec}(\mu,\nu)=\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$ satisfies $\mu$, with $\mu_{\mathrm{in}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{in}}(H))$, and (ii) $\mathrm{dec}(\neg\mu,\nu)=\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\neg\mu$, with $\mu_{\mathrm{out}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{out}}(H))$.
Proof:
(i) Let $\mu_{\mathrm{in}}:=(x\in B_{\mathrm{in}}(H))$. By construction, it holds that $B_{\mathrm{in}}(H)\equiv\mathrm{Proj}_{x_{1}}(B_{\mathrm{in}}(H))\times\cdots\times\mathrm{Proj}_{x_{\nu}}(B_{\mathrm{in}}(H))$, implying that $\mu_{\mathrm{in}}\equiv\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$. Since, $B_{\mathrm{in}}(H)\subseteq H$, $\mu_{\mathrm{in}}$ satisfies $\mu$, which leads to $\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$ satisfying $\mu$ as needed. (ii) Let $\mu_{\mathrm{out}}:=(x\in B_{\mathrm{out}}(H))$. Similarly to (i), it holds that $\mu_{\mathrm{out}}\equiv\wedge_{i=1}^{\nu}\mu_{\mathrm{out}}^{i}$, since $B_{\mathrm{out}}(H)\equiv\mathrm{Proj}_{x_{1}}(B_{\mathrm{out}}(H))\times\cdots\times\mathrm{Proj}_{x_{\nu}}(B_{\mathrm{out}}(H))$, by construction. Then, we have $\neg\mu_{\mathrm{out}}\equiv\lor_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$. Since $\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\lor_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ and $\neg\mu_{\mathrm{out}}$ satisfies $\neg\mu$, we have that $\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\neg\mu$, completing the proof. ∎
Example 2: Consider formulas $\phi=\square_{[a,b]}\mu$ and $\psi=\square_{[a,b]}\neg\mu$, where $\mu:=(h(x)\geq\bm{0})$ is a predicate, with $h(x)=\begin{bmatrix}c_{1}&c_{2}\end{bmatrix}x+\bm{1}_{4}$, $c_{1}=\begin{bmatrix}-\bm{1}_{2}^{\intercal}&\bm{1}_{2}^{\intercal}\end{bmatrix}^{\intercal}$, $c_{2}=\begin{bmatrix}1&-1&1&-1\end{bmatrix}^{\intercal}$. We wish to convert $\phi$ and $\psi$ into conjunctive forms of two conjuncts (subformulas). First, we define the polytope $H=\{x\mid h(x)\geq\bm{0}\}$, and construct orthotopes $B_{\mathrm{in}}(H)=\{x\mid b_{\mathrm{in}}(x)\geq\bm{0}\}$ and $B_{\mathrm{out}}(H)=\{x\mid b_{\mathrm{out}}(x)\geq\bm{0}\}$, with $b_{\mathrm{in}}(x)=2\begin{bmatrix}\bm{I}_{2}&-\bm{I}_{2}\end{bmatrix}^{\intercal}x+\bm{1}_{4}$ and $b_{\mathrm{out}}(x)=\begin{bmatrix}\bm{I}_{2}&-\bm{I}_{2}\end{bmatrix}^{\intercal}x+\bm{1}_{4}$, respectively, as in Definition 4. Polytope $H$ and orthotopes $B_{\mathrm{in}}(H)$, $B_{\mathrm{out}}(H)$ are depicted in Figure 4. Then, using Lemma 5, we define predicates $\mu_{\mathrm{in}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{in}}(H))=(-0.5\leq x_{i}\leq 0.5)$, $\mu_{\mathrm{out}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{out}}(H))=(-1\leq x_{i}\leq 1)$, with $i\in\{1,2\}$, allowing us to decompose $\mu$ and $\neg\mu$ as $\mathrm{dec}(\mu,2)=\mu_{\mathrm{in}}^{1}\wedge\mu_{\mathrm{in}}^{2}$ and $\mathrm{dec}(\neg\mu,2)=\neg\mu_{\mathrm{out}}^{1}\wedge\neg\mu_{\mathrm{out}}^{2}$. Finally, $\mathrm{dec}(\phi,2)=\phi^{1}\wedge\phi^{2}$ and $\mathrm{dec}(\psi,2)=\psi^{1}\wedge\psi^{2}$, where $\phi^{i}=\square_{[a,b]}\mu_{\mathrm{in}}^{i}$ and $\psi^{i}=\square_{[a,b]}\neg\mu_{\mathrm{out}}^{i}$, with $i\in\{1,2\}$, are valid decompositions of $\phi$ and $\psi$, respectively.
[…]
Figure 4: Polytopes $H$, $B_{\mathrm{in}}(H)$, $B_{\mathrm{out}}(H)$, for Example 1.
IV-B Heuristic Agent-Specification Filtering
STL is equipped with robustness metrics for assessing the robust satisfaction of a formula [18]. A scalar-valued function $\rho^{\phi}(\bm{x},k)$ of a signal $\bm{x}$ and time $k$ indicates how robustly a signal $\bm{x}$ satisfies a formula $\phi$ at time $k$. The robustness function is defined recursively in [18, Definition 3].
Let $\phi_{k}$ be a newly assigned STL specification at time $k$, and assume that $\mathrm{dec}(\phi_{k},\nu_{k})=\wedge_{j=1}^{\nu_{k}}\phi_{k}^{j}$. In the following, we assume that the total number of agents, denoted by $|\mathcal{I}|$, satisfies $|\mathcal{I}|>\nu_{k}$, otherwise, the following filtering procedure is ignored. Although, $\phi_{k}$ is split into $\nu_{k}$ agent-level sub-specifications, the assignment of $\phi_{k}^{j}$ to a specific agent $i\in\mathcal{I}$ is not determined by $\mathrm{dec}(\phi_{k},\nu_{k})$. In fact, each sub-specification, $\phi_{k}^{j}$, may be assigned to any agent $i\in\mathcal{I}$ resulting in $\nu_{k}|\mathcal{I}|$ agent-specification pairs. To reduce the number of possible pairs, we utilize the following heuristic approach: Based on the quantitative semantics of STL introduced earlier, we compute the robustness function of $\phi_{k}^{j}$ over the $i$th agent’s trajectory. This metric indicates roughly the distance between the trajectory of agent $i$ and the space of signals that satisfy $\phi_{k}^{j}$ [25], thereby facilitating a preliminary agent-specification assessment. We approximate the $i$th agent’s trajectory by $\bm{z}^{i,k-1}=\{z_{0}^{i,k-1},z_{1}^{i,k-1},\ldots\}$, $i\in\mathcal{I}$, that is, the signal associated with the (deterministic) nominal dynamics of agent $i$, (4a), computed at time $k-1$, which is available at time $k$. Specifically, for each $j\in\{1,\ldots,\nu_{k}\}$, and $i\in\mathcal{I}$, we efficiently calculate $\rho^{\phi_{k}^{j}}(\bm{z}^{i,k-1},k)$, which requires negligible computational cost. Subsequently, for each $\phi_{k}^{j}$, we collect the largest $\nu_{k}$ values of $\rho^{\phi_{k}^{j}}(\bm{z}^{i,k-1},k)$ among the $|\mathcal{I}|$ options, resulting in $(\nu_{k})^{2}$ agent-specification pairs. Should $v_{k}<<|\mathcal{I}|$, the result is a significant reduction.
IV-C Task Allocation Algorithm
Let $\phi$ be a specification which can be decomposed into $\nu$ sub-specifications. After filtering agent-specification pairs and attempting optimization problem (9) with the remainder, a list $\mathbb{L}\subset\mathcal{I}\times\{1,\cdots,\nu\}$ of feasible agent-specification pairs is generated. Each feasible agent-specification pair has a local risk value (10) denoted by $r^{i}_{\phi_{j},mpc}$, with $i\in\mathcal{I}$ and $j\in\{1,\cdots,\nu\}$. Our objective is to auction sub-specifications to agents based on the local risk values, obtaining a definitive assignment in the process. Here, each agent can only receive one specification, and all specifications must be allocated.
To determine an agent-specification assignment, we solve the following binary optimization problem, given by
$\min_{\bm{\lambda}}\ \sum_{(i,j)\in\mathbb{L}}$ $\lambda_{i,j}r^{i}_{\phi_{j},mpc},$ (12a)
[…]
$\sum_{j\in\{1,\cdots,\nu\}}\lambda_{i,j}$ $\leq 1,\ \forall i\in\mathcal{I},$ (12c)
$\lambda_{i,j}$ $=0,\ \ \ \ \ \ \forall(i,j)\notin\mathbb{L},$ (12d)
$\lambda_{i,j}$$\in\{0,1\},\ \forall(i,j)\in\mathbb{L},$ (12e)
where $\bm{\lambda}=\{\lambda_{i,j}\mid(i,j)\in\mathbb{L}\}$. Here, constraint (12b) entails that each specification must be assigned to exactly one agent; constraint (12c) entails that each agent can at most receive one specification; and $\lambda_{i,j}=1$ implies agent $i$ and sub-specification $j$ are chosen as an agent-specification pair. The cost minimizes the sum of the local risk values, thereby ensuring the best probability satisfaction of specification $\phi$. Should no solution be found, specification $\phi$ will be rejected.
IV-D Implementation & Theoretical Analysis
To enact the proposed method, we consider Algorithm 1, which demonstrates multi-agent task allocation and control strategy updates for each agent at every time step. For a more detailed algorithm on single-agent control synthesis, we defer to [6] and the algorithm therein.
Algorithm 1 Task Allocation and Control Strategy Update
1:Given: System (1) initialized at $x^{i}(0)$ for all $i\in\mathcal{I}$
2:Get normalized dynamics of (1) for all $i\in\mathcal{I}$,
3:for $k\in\{0,\cdots,N-1\}$ do
4:     if no new spec. then set $\mathbb{L}_{opt}=\emptyset$ go to line 16
5:     Task Allocation:
6:     ($\psi$, $r_{\psi,\max}$) $\leftarrow$ load new specification
7:     Decompose into ($\psi^{j}$, $r_{\psi^{j},\max}$), $j\in\{1,\cdots,\nu\}$
8:     Filter pairs to obtain $\mathbb{L}\subset\mathcal{I}\times\{1,\cdots,\nu\}$
9:     for $(i,j)\in\mathbb{L}$ do
10:         Solve (9), get $r^{i}_{\psi^{j},mpc}$ and control strategy $\bm{f}^{(i,j)}$
11:         if no solution then $\mathbb{L}=\mathbb{L}\setminus\{(i,j)\}$
12:     end for
13:     Solve (12) to get definitive assignment list $\mathbb{L}_{F}$
14:     Control Strategy Update:
15:     $\forall(i,j)\in\mathbb{L}_{opt}$, set $\bm{f}^{i,k}=\bm{f}^{(i,j)}$leftarrow update
16:     for $i\in\mathcal{I}\setminus\mathrm{Proj}_{\mathcal{I}}(\mathbb{L}_{F})$ do
17:         Solve (9) to get control strategy $\bm{f}^{i}_{meas}$
18:         if solution then set $\bm{f}^{i,k}=\bm{f}^{i}_{meas}$leftarrow update
19:         else set $\bm{f}^{i,k}=\bm{f}^{i,k-1}$leftarrow no update
20:     end for
21:     Implement $\bm{f}^{\mathcal{I},k}$ and measure $x^{i}(k+1),\ \forall i\in\mathcal{I}$
22:end for
Let us consider the recursive feasibility of updating a control strategy for each agent at any time.
Theorem 6 (Recursive feasibility)
If there exists a feasible control strategy $\bm{f}^{\mathcal{I},k}$, at time $k$, then there exists a feasible control strategy $\bm{f}^{\mathcal{I},k+1}$ at time $k+1$.
Proof:
It is sufficient to prove that the existence of $\bm{f}^{i,k}$ implies the existence of $\bm{f}^{i,k+1}$ for each agent $i\in\mathcal{I}$. Notice that should an agent either accept a new sub-specification or measurement only, an updated control strategy is obtained, respectively, as $\bm{f}^{i,k+1}=\bm{f}^{(i,j)}$ or $\bm{f}^{i,k+1}=\bm{f}^{i}_{meas}$. Should neither be available, the previous control strategy is again implemented, $\bm{f}^{i,k+1}=\bm{f}^{i,k}$, completing the proof. ∎
For each specifications $\phi$ accepted at time $k$, by subgroup $I\subseteq\mathcal{I}$, we need that $\mathbb{P}_{\bm{f}}(\phi,k)\geq 1-r_{\phi,\max}$. Here, it is sufficient to show that $\forall i\in I$, $\mathbb{P}_{\bm{f}}(\phi^{i},k)\geq 1-r_{\phi^{i},\max}$. The latter, in relation to optimization problem (9), is discussed within open-loop and closed-loop implementation in our previous work [6]. In said paper, we established the validity of open-loop implementation and suggested that closed-loop implementation follows naturally from the convex unimodality of the disturbance.
Remark 1
Note that the local risk value $r_{\phi^{j},mpc}$ compared to the average maximal risk $r_{\phi^{j},\max}$ is conservative, $r_{\phi^{j},mpc}\geq r_{\phi^{j},\max}$, due to a union-bound argument used in proving Prop. 3. Also, the obtained task allocation may not always be optimal or feasible due to our heuristic choice of filtering, even if an allocation exists.
V Case Study
[…]
In addition to global and local specifications investigated in this paper, the scenario also allows for specifications explicitly assigned to individual agents, such as routing a specific bus to a designated region. The overall dynamically allocated specifications in this scenario are listed as follows.
  • When initialized, each bus X starts from T-X and is assigned a safety specification $\varphi_{(0,X)}\!:=\!\square_{[0,N]}(\mu_{\mathrm{BOX}}\wedge\lnot\mu_{\mathrm{B}\textbf{-}1}\wedge\lnot\mu_{\mathrm{B}\textbf{-}2})$.
  • At time $k=2$, a global task is given to the buses, asking them to transport tourists from GP-I, II, IV to ULP within 15 time steps. We decompose the global specification into three local specifications given by:
    $\varphi_{(2,\mathrm{I})}\!:=\!\lozenge_{[2,9]}\mu_{\mathrm{GP}\textbf{-}\mathrm{I}}\!\wedge\!\square_{[2,9]}(\mu_{\mathrm{GP}\textbf{-}\mathrm{I}}\!\rightarrow\!\lozenge_{[0,7]}\mu_{\mathrm{ULP}})$,
    $\varphi_{(2,\mathrm{II})}\!:=\!\lozenge_{[3,10]}\mu_{\mathrm{GP}\textbf{-}\mathrm{II}}\!\wedge\!\square_{[3,10]}(\mu_{\mathrm{GP}\textbf{-}\mathrm{II}}\!\rightarrow\!\lozenge_{[0,7]}\mu_{\mathrm{ULP}})$,
    $\varphi_{(2,\mathrm{IV})}\!:=\!\lozenge_{[4,11]}\mu_{\mathrm{GP}\textbf{-}\mathrm{IV}}\!\wedge\!\square_{[4,11]}(\mu_{\mathrm{GP}\textbf{-}\mathrm{IV}}\!\rightarrow\!\lozenge_{[0,7]}\mu_{\mathrm{ULP}})$.
  • At time $k=8$, a global task is given to the buses, asking them to transport tourists from GP-II, III to ULP within 15 time steps. We decompose the global specification into two local specifications given by:
    $\varphi_{(8,\mathrm{III})}\!:=\!\lozenge_{[8,15]}\mu_{\mathrm{GP}\textbf{-}\mathrm{III}}\!\wedge\!\square_{[8,15]}(\mu_{\mathrm{GP}\textbf{-}\mathrm{III}}\!\rightarrow\!\lozenge_{[0,7]}\mu_{\mathrm{ULP}})$,
    $\varphi_{(8,\mathrm{II})}\!:=\!\lozenge_{[9,16]}\mu_{\mathrm{GP}\textbf{-}\mathrm{II}}\!\wedge\!\square_{[9,16]}(\mu_{\mathrm{GP}\textbf{-}\mathrm{II}}\!\rightarrow\!\lozenge_{[0,7]}\mu_{\mathrm{ULP}})$.
  • At time $k=15$, each bus $X$ is assigned with a ‘back to terminal’ specification $\varphi_{(15,X)}\!:=\!\lozenge_{[15,23]}\square_{[0,2]}\mu_{\mathrm{T}\textbf{-}\mathrm{X}}$.
All four buses have the same dynamic models described by system (1) with $A\!=\!B\!=\!I_{2}$ and $w^{i}(k)\sim\mathcal{N}(0,\sigma I_{2})$ for all $i\in\mathcal{I}$, with $\sigma=0.001$. The states and the control inputs are their positions and velocities, respectively. To incorporate the heterogeneity of practical systems, we assume that the buses have different speed limits, i.e., 444\,4m/s, 555\,5m/s, 666\,6m/s, and 777\,7m/s for buses A to D, respectively. We consider a stabilizing feedback gain $K\approx-0.618I_{2}$ for Eq. (4b), a maximal risk $r_{*,\max}=0.5$ for any global specification or explicit specification ($k=0$, $k=15$), individual maximal risk based on Eq. (11) for any local specification ($k=2$, $k=8$) and a cost $J(\bm{z},\bm{v},\bm{r})=u^{T}Ru+\sum_{i=1}^{N}r(i)$ for Eq. (9a) with $R=0.001I_{2}$. A coordinate transformation ensures all PRS are spherical and the dynamics are normalized.
[…]
After decomposing, filtering, control synthesis, and auctioning, we obtain the results as seen in Figure 5.
Figure 5: The bus trajectories.
The logging information of the experiment indicates that: 1) at time $k=2$, buses A, B, and D are assigned with $\varphi_{(2,\mathrm{II})}$, $\varphi_{(2,\mathrm{IV})}$, and $\varphi_{(2,\mathrm{I})}$, respectively; 2) at time $k=8$, buses A and D are assigned with $\varphi_{(8,\mathrm{II})}$ and $\varphi_{(8,\mathrm{III})}$, respectively; 3) at time $k=15$, buses B, C, and D have accepted the return specification, while bus A has rejected it. Bus C is not assigned any task throughout the experiment and, as a result, stays at its terminal. Since buses A and D accept additional tasks at $k=8$, our method allows agents to adjust their behaviours according to newly assigned tasks. As mentioned, bus A has rejected the return task at time $k=15$. This is due to the incapability to move long distances considering its small speed limit. Additionally, all buses are confined within the ‘BOX’ region and kept outside the buildings, implying the compatibility of our method to explicitly assigned specifications. Part of the trajectory of bus D intersects with B-2 due to naive linear interpolation. In practice, critical safety can be guaranteed using nonlinear interpolation which may generate curved trajectories navigating around the building.
Remark 2
In this experiment, the computation of control inputs by solving problem (9) is distributed to each agent, leading to a computational complexity linear to the number of assigned specifications. In this sense, our proposed method allows for a comparable computational load of each agent to a conventional single-agent case, allowing it to be implemented in practical scenarios.
VI Conclusion
This paper considered real-time task allocation and control synthesis for stochastic heterogeneous linear multi-agent systems with probabilistic satisfaction of signal temporal logic specifications. The proposed method first decomposed the specification onto the single-agent level, to next use robustness of STL to filter the most promising agent-specification pairs. Subsequently, tube-based model predictive control was used to synthesize a control strategy and determine the local risk value for each remaining agent-specification pair. Finally, an auctioning algorithm based on the local risk values determined the optimal assignment. The proposed method was demonstrated through a multi-bus scenario.
In future work, we will consider resolving conflict due to the rejection of specifications and consider a fully centralized control synthesis. Additionally, improvements upon the heuristic approach and the conservativeness due to union bound argument will be considered.